A monopolist has a cost curve c(q) = q^2-12q+8 and faces an inverse demand curve p(q) = 80-20q. Find the monopolist price and quantity, (p,q).
Answer #1

Date Page - C = 9 - 129 +8 Me – de 09 - d (92) - 12.d dq (g) + d (8) I da da da = 29 - 12 - P: 80 - 209 TR 2 800 - 209 MR = 8

P=36.2

Q=2.19